# Counterfactual Explanations with Probabilistic Guarantees on their Robustness to Model Change

### v1

@@ Abstract
Machine Learning, Counterfactual Explanations, Robustness, Explainable AI
@@ 1 Introduction
Counterfactual explanations (counterfactuals, CFEs) are one of the most popular forms of explaining decisions made by complex, black-box machine learning (ML) algorithms. Briefly, a counterfactual explanation of a decision $y$ made for input $x$ is an instance $x^{cf}$ that is very similar to $x$ but produces a different, more desirable prediction $y^{\prime}\neq y$. Since CFEs can be interpreted as an answer to the question: “given the decision $y$ taken for input $x$, how should $x$ be changed to produce the alternative decision $y^{\prime}$?”, they offer actionable feedback to the user. This is appreciated by stakeholders in various application areas such as supporting loan decisions (Wachter et al., 2017), job recruitment (Pearl et al., 2016), medicine (Mertes et al., 2022), and many others (Guidotti, 2022).
Even though the basic definition of a counterfactual (Wachter et al., 2017) specifies only two basic properties: validity (ensuring the desired classification $y^{\prime}$) and proximity (small distance between $x$ and $x^{cf}$), many additional properties are useful from both the application (Keane et al., 2021) and user (Förster et al., 2020; Keane et al., 2021) point of view. These properties include sparsity (modifying values of only few features), actionability (realistic feature changes), plausibility (proximity to the data distribution), and many others. Even tough numerous methods for generating counterfactuals with different properties have been proposed (Guidotti, 2022), almost all of them deal with a static problem setting and overlook the problem of counterfactual robustness to model change.
Since a counterfactual is intended to deliver actionable feedback to the user, it must remain valid for the period of time necessary to undertake changes. Note that a CFE is generated for a fixed model, however, many applications are inherently dynamic, and the model is changing over time. There are many possible causes of such model change, some of which include obtaining new training data, adjusting hyperparameters or model architecture, and even the need to remove some of the training data due to data expiration policies or privacy laws (someone may request their personal data to be removed (Ginart et al., 2019)). In such scenario, it is important to preserve the validity of the counterfactual for the newly retrained model, so that the user will still get the desired decision while acting on the recourse offered to them before the model change.
@@
This highlights the need for a fresh look at the properties of counterfactuals in the context of changing environments, and in particular leads us to the concept of robustness to model change. This challenge has been considered from various perspectives, including robustness to input perturbations (Artelt et al., 2021; Ehyaei et al., 2023), to model changes (Upadhyay et al., 2021; Dutta et al., 2022; Jiang et al., 2023) or imperfect realizations of recommendations (Pawelczyk et al., 2022; Guyomard et al., 2023; Maragno et al., 2024). In this paper, we focus on the robustness to moderate model changes, where we expect the CFE to remain valid over time with regard to the class indicated before retraining. This topic has not been thoroughly explored yet, and the existing introduced methods exhibit significant limitations. Some of these methods rely on impractical assumptions, while others are restricted to specific models and often require substantial human effort to tune their non-interpretable hyperparameters. Most importantly, they typically provide no statistical guarantees of robustness and do not specify against which model changes the counterfactual is robust.
@@
Fig. 1 illustrates the essence of BetaRCE. Initially, we obtain the base counterfactual (depicted as a green square) using a base explanation method for a given input example (green triangle). Subsequently, the method moves this base counterfactual to a safer data region, considering the probability of change in classification within that area. Here, $\delta$ represents the lower bound of the probability of a counterfactual preserving its original class. Higher $\delta$ values correspond to a higher degree of robustness, which typically comes with a more substantial deviation from the base counterfactual.
To validate our method, we conducted experiments that confirmed the introduced probabilistic bounds hold in practice for various types of model changes. Furthermore, the experiments showed that counterfactuals generated by BetaRCE are closer to the base CFE and better preserve their properties compared to previous methods aimed at improving robustness.
@@ 2 Related Work
Counterfactual explanations are a widely explored topic in the contemporary ML literature. Some well-known representatives of such methods (with an indication of the counterfactual properties they try to take into account) include Wachter (Wachter et al., 2017) – closest to the input instance, GrowingSpheres (Laugel et al., 2018) – closest and sparse, FACE (Poyiadzi et al., 2020) – located in dense regions, Dice (Mothilal et al., 2020) – multiple closest, sparse, plausible and diverse counterfactuals. For a comprehensive survey, see e.g., (Guidotti, 2022; Verma et al., 2020). These methods generate counterfactuals with various properties considered useful from the user’s perspective. However, Stepka et al. (2024) used multi-criteria analysis to demonstrate that none of the methods excels in all evaluation measures simultaneously.
It should be noted that these methods usually do not consider the issue of robustness, as this challenge has been noticed relatively recently (Mishra et al., 2023). Nonetheless, the issue, due to its practical importance, has already been examined from several perspectives. In the following, we briefly outline the three main definitions of CFE’s robustness:
Robustness to input perturbations requires that similar input examples should yield similar CFEs, i.e. the difference between them should be minimized. This idea is closely related to the notion of individual fairness, where two similar individuals are expected to receive similar treatment (Artelt et al., 2021; Ehyaei et al., 2023; Artelt & Hammer, 2022).
Robustness to imperfect recourse refers to a situation when the user acting on the counterfactual recommendations slightly fails to meet the exact proposed changes. It is argued, that actionable recourse should allow slight imperfections in the realization of the recommendation (Pawelczyk et al., 2022). Proposed methods (Maragno et al., 2024; Pawelczyk et al., 2022; Guyomard et al., 2023; Virgolin & Fracaros, 2023) generally aim to ensure that the class distribution in some neighbourhood of the CFE is relatively pure in order to account for the mentioned imperfect execution of the recommendation.
@@
Up to this point, some methods attempting to solve this type of robustness have been proposed, e.g., see this recent survey (Jiang et al., 2024). However, most of the introduced methods either do not quantify the extent of achieved robustness, or if they do, they are inherently model-specific, constraining the versatility of such approach. On top of that, they also do not enable decision-makers to directly specify their expectations regarding the desired level of robustness. Below, we briefly describe a selection of existing methods employing varied paradigms of ensuring robustness.
Upadhyay et al. (2021) introduced ROAR, an end-to-end method that generates robust CFEs using a training procedure with a custom adversarial objective that optimises for the worst perturbation of the locally approximated decision boundary. Nguyen et al. (2022) proposed RBR, a Bayesian-inspired method that models the data distribution with Gaussian kernels and accounts for data perturbations. Both previously mentioned methods only account for model changes due to data shifts and require significant effort in hyperparameter tuning. In (Ferrario & Loi, 2022), robustness is achieved by modifying the training procedure of the model itself through counterfactual data augmentation, which may affect model performance as well as applicability to various practical settings.
The idea of post-hoc CFE generation was introduced with the RobX method(Dutta et al., 2022), which was designed for non-differentiable tree-based models, but is in principle applicable to any black-box model. RobX explores the local neighbourhood of a counterfactual by querying the underlying decision model over synthetic data points, thereby assessing local class variability. However, RobX has notable limitations. It relies heavily on the assumption that the decision model is well calibrated. It also requires dataset-specific fine-tuning of the hyperparameters, which significantly affects the performance of the method. These hyperparameters include $\tau$ (the threshold for the counterfactual stability metric) and variance (representing the size of the local neighbourhood used for sampling). In scenarios where the model can change, it is uncertain whether RobX quantifies the true local variability with high fidelity.
@@ 3.1 Defining robustness
Consider a binary classification problem where a machine learning model $M:\mathcal{X}\rightarrow\mathcal{Y}$ assigns a binary label $y\in\mathcal{Y}=\{0,1\}$ to each instance $x\sim\mathcal{X}$. Given an input instance $x^{orig}$ and a prediction $M(x^{orig})$, a counterfactual explanation $x^{cf}$ is an instance similar [^1] Different notions of similarity between the counterfactual and the original instance have been used in methods presented in the literature. to $x^{orig}$ for which the model reaches the opposite decision, i.e. $M(x^{orig})\neq M(x^{cf})$.
@@
Definition 1 (Robust counterfactual).
@@
Definition 2 (Space of admissible models).
@@
In the related literature, model changes are defined as perturbations of model parameters (Dutta et al., 2022; Xu & Wu, 2024; Jiang et al., 2023). Here, we use a more general definition of a model change, which is defined by the user, and can include changes in random seeds, model hyperparameters, or changes in the training data set. Since the space of admissible models is potentially infinite, we relax the notion of counterfactual’s robustness to only some (sufficiently high) proportion of possible admissible models.
Definition 3 ($\delta$-robust counterfactual).
@@ 3.2 Estimating robustness
We adopted a Bayesian perspective on the estimation of $\delta$ to account for the estimation error. Since robustness to a given model change is a binary variable following a Bernoulli distribution, we used the default prior for binary data, Beta distribution, to model the confidence of the estimate $\hat{\delta}$. Recall that Beta is the conjugate prior of the Bernoulli distribution (Gelman et al., 2013), which allows for much simpler computations. The adoption of the concept of credible interval, specifically its lower bound, leads to the following definition, which accounts for the estimation error.
Definition 4 ($(\delta,\alpha)$-robust counterfactual).
@@
In simple words, $(\delta,\alpha)$-robust CFE at $\alpha$ confidence level has the probability of being robust of at least $\delta$. This is reminiscent of the classic definition of PAC learning (Haussler & Warmuth, 1993), where we obtain a counterfactual that is likely to be approximately ($\delta$) robust.
A practical procedure for verifying $(\delta,\alpha)$-robustness for a given counterfactual $x^{cf}$ can be implemented using various statistical techniques, but in this work, we show that using simple bootstrap estimation of the parameter $\hat{\delta}$ is enough to obtain useful results. More concretely, we sample $k$ estimators from the admissible model space $M^{\prime}\sim\mathcal{M}_{M}$ and apply them to obtain a set of predictions for the counterfactual $M^{\prime}(x^{cf})$. The outcomes compared to the desired class $M(x^{cf})$ are used to update the noninformative Jeffreys prior (Gelman et al., 2013) for $\hat{\delta}$ following Beta distribution. Finally, the verification of the condition described in Def. 4 can be validated via checking the quantile of a posteriori Beta distribution.
@@
Note that although the above procedure requires sampling and thus retraining of multiple models, the same small sample of models can be used to check all counterfactuals, greatly reducing the computational requirements as all the models can be trained and stored beforehand. Similarly, the use of the conjugate Beta distribution allows for very effective Bayesian updating of the posterior distribution given the observed data. The pseudocode of the verification procedure is given in Alg. 1.
Theorem 1.
@@
The proof can be found in App. B.
Algorithm 1 Bootstrap verification of $(\delta,\alpha)$-robustness
  Input
  $x^{cf},y^{cf}$ counterfactual explanation and its desired class
  $\mathcal{M}$ space of admissible models
  $k$ number of estimators
  Procedure
  $a,b\leftarrow(0.5,0.5)$Initialize the parameters of a priori beta distribution (noninformative Jeffreys)}
  for $i\in 1..k$ do
     $M^{\prime}\leftarrow\text{a sample from }\mathcal{M}_{M}$Update posterior distribution}
     if $M^{\prime}(x^{cf})==y^{cf}$ then
        $a\leftarrow a+1$
     else
        $b\leftarrow b+1$
     end if
  end for
  return  $F^{-1}_{Beta(a,b)}(1-\alpha)\geq\delta$Check the condition for counterfactual robustness (Def. 4)}
@@ 3.3 BetaRCE - a post-hoc method for making CFEs $(\delta,\alpha)$-robust
Multiple counterfactual construction methods have been proposed in the literature, and according to various studies, the selection of the most appropriate counterfactual strongly depends on user preferences (Stepka et al., 2024). Therefore, we present BetaRCE , a post-hoc approach that generates a $(\delta,\alpha)$-robust counterfactual by making a small perturbation to the counterfactual $x^{cf}$ constructed by a method selected by the user (to meet his expectations regarding the selected evaluation measures).
@@
$x^{*}=\operatorname*{argmin}_{x^{\prime}\in\mathcal{X}}{d(x^{cf},x^{\prime})\quad\text{s.t.}\quad valid(x^{\prime},x^{orig})\land robust(x^{\prime})}$ (6)
@@
The above objective formulation (Eq. 6) is non-convex and non-differentiable, so any zero-order optimization algorithm could be used. Here we chose GrowingSpheres (Laugel et al., 2018) as it is a simple and fast optimization method originally designed to find adversarial examples that closely resemble the original input while inducing sparse feature changes. It is therefore well suited to finding a robust counterfactual that closely resembles the original one, especially since it can directly optimize our objective Eq. 6. However, we acknowledge that many other zero-order optimization methods could be used to optimize that objective.
GrowingSpheres performs the optimization in two main steps, each involving the generation of examples uniformly distributed in a sphere constructed around the given instance $x^{cf}$. In the first step, the method searches for the largest radius of a sphere containing only non-robust or invalid instances. The optimization is iterative, halving the original large radius of a sphere until a lower bound on the distance from $x^{cf}$ to both valid and robust counterfactuals is established. In the second step, the method repeats the optimization process in the opposite direction, generating random instances in the iteratively growing sphere, but not closer than the estimated lower bound. The optimization process ends when the first valid and robust counterfactual is found. GrowingSpheres has two hyperparameters: $\eta$, which denotes the initial sphere radius, and $n$, representing the number of sampled instances from a sphere. We set these parameters to the default values from the original paper. The pseudocode of BetaRCE is depicted in Alg. 2. The function $Sphere$ is a sampling procedure employed in GrowingSpheres, with arguments respectively denoting the center to sample around, the shorter radius, and the longer radius (sampling occurs between these two radii).
Remark 1.
@@
  Input
  $x^{orig},y^{orig}$ input instance and its class
  $\delta$, $\alpha$, $\mathcal{M}_{M}$, $k$ BetaRCE hyperparameters
  $\eta$, $n$ GrowingSpheres hyperparameters
  Procedure
  $x^{cf}\leftarrow BaseCFE(M,x^{orig},y^{orig})$Constructing the original counterfactual}
  {Estimating lower bound on distance to robust and valid CFEs}
  $Z\sim_{n}Sphere(x^{cf},0,\eta)$Uniformly sample $n$ candidates}
  while $\{x^{\prime}\in Z:valid(x^{\prime})\land robust(x^{\prime})\}\neq\emptyset$ do
     $\eta\leftarrow 0.5\eta$
     $Z\sim_{n}Sphere(x^{cf},0,\eta)$
  end while
  {Growing spheres to find robust and vaild CFEs}
  $a_{0}\leftarrow\eta$, $a_{1}\leftarrow 2\eta$
  $Z\sim_{n}Sphere(x^{cf},a_{0},a_{1})$
  while $\{x^{\prime}\in Z:valid(x^{\prime})\land robust(x^{\prime})\}==\emptyset$ do
     $a_{0}\leftarrow a_{1}$,
     $a_{1}\leftarrow a_{1}+\eta$
     $Z\sim_{n}Sphere(x^{cf},a_{0},a_{1})$
  end while
  return  $\min_{x^{\prime}\in Z:valid(x^{\prime})\land robust(x^{\prime})}d(x^{\prime},x^{cf})$Return robust and valid CFE closest to the orignal CFE}
Figure 2: The average empirical robustness of counterfactuals generated by BetaRCE at $\alpha=95\%$, with GrowingSpheres generating base CFEs. Red and green dashed lines are average lower and upper bounds of estimated $\alpha$ credible intervals. The shaded areas in the back represent standard deviations. Horizontal yellow and black lines show the base robustness of CFEs obtained with Dice and GrowingSpheres without BetaRCE applied. The plots with all four datasets and two different BetaRCE base CFE generation methods (Dice and GrowingSpheres ) are available in App.G.1.
@@
Proof sketch: In Theorem 2 $a+k$ represents the most optimistic situation, where all $k$ estimators agree that the counterfactual is robust, resulting in the most extreme right-skewed beta distribution and the highest attainable $\delta$ with $k$ estimators. Proof and a more detailed analysis of this formula are available in App. C.

### v3

@@ 1 Introduction
Counterfactual explanations (CFEs) are one of the most popular forms of explaining decisions made by complex, black-box machine learning (ML) algorithms. Briefly, a counterfactual explanation of a decision $y$ made for input $x$ is an instance $x^{cf}$ that is very similar to $x$ but produces a different, more desirable prediction $y^{\prime}\neq y$. Since CFEs can be interpreted as an answer to the question: “given the decision $y$ taken for input $x$, how should $x$ be changed to produce the alternative decision $y^{\prime}$?”, they offer actionable feedback to the user. Over the years, it has been appreciated by stakeholders in various application areas such as supporting loan decisions [1], job recruitment [2], medicine [3], and many others [4].
Even though the basic definition of a counterfactual [1] specifies only two basic properties: validity (ensuring the desired classification $y^{\prime}$) and proximity (small distance between $x$ and $x^{cf}$), many additional properties are useful from both the application [5] and user [6, 5] point of view. These properties include sparsity (modifying values of only few features), actionability (realistic feature changes), plausibility (proximity to the data distribution), and many others. Numerous methods for generating counterfactuals with different properties have been proposed [4], yet almost all of them deal with static problem settings, overlooking the counterfactuals’ robustness to model change.
Since counterfactuals are intended to deliver actionable feedback, they must remain valid for the period of time that is necessary for the end-user to act on the proposed changes. Note that CFEs are generated for a fixed model, yet many applications are inherently dynamic with the model changing over time. There are many possible causes incurring a model change, some of which include obtaining new training data, adjusting hyperparameters or model architecture, and having to remove some training data due to data expiration policies or privacy laws (someone may request their personal data to be removed [7]). In such scenario, it is important to preserve the validity of the counterfactual for the newly retrained model, so that the user will still get the desired decision while acting on the recourse offered to them before the model change.
@@
This highlights the need for a fresh look at the properties of counterfactuals in the context of changing environments, and in particular, leads us to the concept of robustness to model change. This challenge has been considered from various perspectives, including robustness to input perturbations [8, 9], to model changes [10, 11, 12] or imperfect realizations of recommendations [13, 14, 15]. In this paper, we focus on the robustness to moderate model changes, where we expect the CFE to remain valid over time with regard to the class indicated before retraining. This topic has not been thoroughly explored yet, and the existing introduced methods exhibit significant limitations. Some of these methods rely on impractical assumptions, while others are restricted to specific models and often require substantial human effort to tune their non-interpretable hyperparameters. Most importantly, they typically provide no statistical guarantees of robustness and do not specify against which model changes the counterfactual is robust.
@@
Fig. 1 illustrates the essence of BetaRCE. Initially, we obtain the base CFE using a base explanation method for a given input example. Subsequently, we move this base CFE to a safer data region, i.e., an area with a low probability of change in classification. Here, $\delta$ represents the lower bound of the probability of a CFE preserving its original class. Higher $\delta$ values correspond to a higher degree of robustness, which often comes with a more substantial deviation from the base CFE. Crucially, in BetaRCE we let the user select the desired $\delta$, giving them the opportunity to decide on the tradeoff between the explanation’s degree of robustness and proximity to the original example.
To validate our method, we conducted experiments that confirmed the introduced probabilistic bounds hold in practice for various types of model changes. Furthermore, the experiments showed that counterfactuals generated by BetaRCE remain close to the base CFE and better preserve their properties compared to previous methods aimed at improving robustness.
@@ 2 Related Work
Counterfactual explanations are a widely explored topic in the contemporary ML literature. Some well-known representatives of such methods (with an indication of the counterfactual properties they try to take into account) include Wachter [1] – closest to the input instance, GrowingSpheres [16] – closest and sparse, FACE [17] – located in dense regions, Dice [18] – multiple closest, sparse, plausible and diverse counterfactuals. Over the years numerous methods were introduced, all suited for various motivations and aiming to attain certain properties that are considered useful from the perspective of the application scenario.
Notably, Stępka et al. [19] demonstrated through multi-criteria analysis that no single method excels across all evaluation measures simultaneously, highlighting the need for CFE methods tailored to specific application requirements. While plenty of these methods have significantly advanced the field, they fall outside the scope of this paper, as we focus specifically on robustness. For comprehensive surveys on counterfactual methods, we direct readers to [4, 20].
Most existing methods, however, do not explicitly address the issue of robustness – a challenge that has only recently gained attention [21]. Nonetheless, due to its practical importance, robustness has already been explored from various perspectives.
In the following, we briefly outline the three primary definitions of counterfactual robustness:
Robustness to input perturbations requires that similar input examples should yield similar CFEs, i.e. the difference between them should be minimized. This idea is closely related to the notion of individual fairness, where two similar individuals are expected to receive similar treatment [8, 9, 22].
Robustness to imperfect recourse refers to a situation when the user acting on the counterfactual recommendations slightly fails to meet the exact proposed changes. It is argued, that actionable recourse should allow slight imperfections in the realization of the recommendation [13]. Proposed methods [15, 13, 14, 23] generally aim to ensure that the class distribution in some neighborhood of the CFE is relatively pure in order to account for the mentioned imperfect execution of the recommendation.
@@
Up to this point, a few methods attempting to solve this type of robustness have been proposed, as covered by a recent survey [24]. However, most of the introduced methods either do not quantify the extent of achieved robustness, or if they do, they are inherently model-specific, constraining the versatility of such approach. On top of that, they also do not enable decision-makers to directly specify their expectations regarding the desired level of robustness. Below, we briefly describe a selection of existing methods employing varied paradigms of ensuring robustness.
Upadhyay et al. [10] introduced ROAR, an end-to-end method that generates robust CFEs using a training procedure with a custom adversarial objective that optimizes for the worst perturbation of the locally approximated decision boundary. Nguyen et al. [25] proposed RBR, a Bayesian-inspired method that models the data distribution with Gaussian kernels and accounts for data perturbations. Both previously mentioned methods only account for model changes due to data shifts and require significant effort in hyperparameter tuning. In [26], robustness is achieved by modifying the training procedure of the model itself through counterfactual data augmentation, which may affect model performance as well as applicability to various practical settings.
The idea of post-hoc CFE generation was introduced with Dutta et al. [11]’s RobX method, designed for non-differentiable tree-based models, but in principle applicable to any black-box model. RobX explores the local neighborhood of a counterfactual by querying the underlying decision model over synthetic data points, thereby assessing local class variability. However, RobX has notable limitations. It relies heavily on the assumption that the decision model is well calibrated. It also requires dataset-specific fine-tuning of the hyperparameters, which significantly affects the performance of the method. These hyperparameters include $\tau$ (the threshold for the counterfactual stability metric) and variance (representing the size of the local neighborhood used for sampling). In scenarios where the model can change, it is uncertain whether RobX quantifies the true local variability with high fidelity.
@@ 3.1 Defining robustness
Consider a binary classification problem where a machine learning model $M:\mathcal{X}\rightarrow\mathcal{Y}$ assigns a binary label $y\in\mathcal{Y}=\{0,1\}$ to each instance $x\sim\mathcal{X}$. Given an input instance $x^{orig}$ and a prediction $M(x^{orig})$, a counterfactual explanation $x^{cf}$ is an instance close to $x^{orig}$ for which the model reaches the opposite decision, i.e. $M(x^{orig})\neq M(x^{cf})$.
@@
Definition 1 (Robust counterfactual)
@@
Definition 2 (Space of admissible models)
@@
In the related literature, model changes are defined as perturbations of model parameters [11, 29, 12]. Here, we use a more general definition of a model change, which is defined by the user, and can include changes in random seeds, model hyperparameters, or changes in the training data set. Since the space of admissible models is potentially infinite, we relax the notion of counterfactual’s robustness to only some (sufficiently high) proportion of possible admissible models.
Definition 3 ($\delta$-robust counterfactual)
@@ 3.2 Estimating robustness
We adopted a Bayesian perspective on the estimation of $\delta$ to account for the estimation error. Since robustness to a given model change is a binary variable following a Bernoulli distribution, we used the default prior for binary data, Beta distribution, to model the confidence of the estimate $\hat{\delta}$. Recall that Beta is the conjugate prior of the Bernoulli distribution [30], which allows for much simpler computations. The adoption of the concept of credible interval, specifically its lower bound, leads to the following definition, which accounts for the estimation error.
Definition 4 ($(\delta,\alpha)$-robust counterfactual)
@@
In simple words, $(\delta,\alpha)$-robust CFE at $\alpha$ confidence level has the probability of being robust of at least $\delta$. This is reminiscent of the classic definition of PAC learning [31], where we obtain a counterfactual that is likely to be approximately ($\delta$) robust.
A practical procedure for verifying $(\delta,\alpha)$-robustness for a given counterfactual $x^{cf}$ can be implemented using various statistical techniques, but in this work, we show that using simple bootstrap estimation of the parameter $\hat{\delta}$ is enough to obtain useful results. More concretely, we sample $k$ estimators from the admissible model space $M^{\prime}\sim\mathcal{M}_{M}$ and apply them to obtain a set of predictions for the counterfactual $M^{\prime}(x^{cf})$. The outcomes compared to the desired class $M(x^{cf})$ are used to update the noninformative Jeffreys prior ([30]) for $\hat{\delta}$ following Beta distribution. Finally, the verification of the condition described in Def. 4 can be validated via checking the quantile of a posteriori Beta distribution.
@@
Note that although the above procedure requires sampling and thus retraining of multiple models, the same small sample of models can be used to check all counterfactuals, greatly reducing the computational requirements as all the models can be trained and stored beforehand. Similarly, the use of the conjugate Beta distribution allows for very effective Bayesian updating of the posterior distribution given the observed data. The pseudocode of the verification procedure is presented in Alg. 1.
Algorithm 1 Bootstrap verification of $(\delta,\alpha)$-robustness
1:Input
2:$x^{cf},y^{cf}$ counterfactual explanation and its desired class
$\mathcal{M}$ space of admissible models
4:$k$ number of estimators
5:Procedure
6:$a,b\leftarrow(0.5,0.5)$triangleright Initialize the parameters of a priori beta distribution (noninformative Jeffreys)
7:for $i\in 1..k$ do
8:     $M^{\prime}\leftarrow\text{a sample from }\mathcal{M}_{M}$
9$\triangleright$ Update posterior distribution
10:     if $M^{\prime}(x^{cf})==y^{cf}$ then
11:         $a\leftarrow a+1$
12:     else
13:         $b\leftarrow b+1$
14:     end if
15:end for
16:return $F^{-1}_{Beta(a,b)}(1-\alpha)\geq\delta$triangleright Check the condition for counterfactual robustness (Def. 4)
Theorem 1
@@
The proof can be found in App. B [^2] The full appendix is available online at: https://ignacystepka.com/projects/betarce.html.
@@ 3.3 BetaRCE - a post-hoc method for making CFEs $(\delta,\alpha)$-robust
Multiple counterfactual construction methods have been proposed in the literature, and according to various studies, the selection of the most appropriate counterfactual strongly depends on user preferences [19]. Therefore, we present BetaRCE , a post-hoc approach that generates a $(\delta,\alpha)$-robust counterfactual by making a small perturbation to the counterfactual $x^{cf}$ constructed by a method selected by the user (to meet his expectations regarding the selected evaluation measures).
@@
$x^{*}=\operatorname*{argmin}_{x^{\prime}\in\mathcal{X}}{d(x^{cf},x^{\prime})~{}~{}\text{s.t.}~{}~{}valid(x^{\prime},x^{orig})\land robust(x^{\prime})}$ (6)
@@
The above objective formulation (Eq. 6) is non-convex and non-differentiable, so any zero-order optimization algorithm could be used. Here we chose GrowingSpheres [16] as it is a simple and fast optimization method originally designed to find adversarial examples that closely resemble the original input while inducing sparse feature changes. It is therefore well suited to finding a robust counterfactual that closely resembles the original one, especially since it can directly optimize our objective Eq. 6. However, we acknowledge that many other zero-order optimization methods could be used to optimize that objective.
GrowingSpheres performs the optimization in two main steps, each involving the generation of examples uniformly distributed in a sphere constructed around the given instance $x^{cf}$. In the first step, the method searches for the largest radius of a sphere containing only non-robust or invalid instances. The optimization is iterative, halving the original large radius of a sphere until a lower bound on the distance from $x^{cf}$ to both valid and robust counterfactuals is established. In the second step, the method repeats the optimization process in the opposite direction, generating random instances in the iteratively growing sphere, but not closer than the estimated lower bound. The optimization process ends when the first valid and robust counterfactual is found. GrowingSpheres has two hyperparameters: $\eta$, which denotes the initial sphere radius, and $n$, representing the number of sampled instances from a sphere. We set these parameters to the default values from the original paper. The pseudocode of BetaRCE is depicted in Alg. 2. The function $Sphere$ is a sampling procedure internally employed in GrowingSpheres, with arguments respectively denoting the center to sample around, the shorter radius, and the longer radius (sampling occurs between these two radii). We provide an intuitive visualization of all steps of the algorithm in App. H.
Remark 1
@@
1:Input
2:$x^{orig},y^{orig}$ input instance and its class
3:$\delta$, $\alpha$, $\mathcal{M}_{M}$, $k$ BetaRCE hyperparameters
4:$\eta$, $n$ GrowingSpheres hyperparameters
5:Procedure
6$\triangleright$ Construction of the base CFE
7:$x^{cf}\leftarrow BaseCFE(M,x^{orig},y^{orig})$
8$\triangleright$ GrowingSpheres warmup: estimate the lower bound on the distance to robust and valid CFEs
9:$Z\sim_{n}Sphere(x^{cf},0,\eta)$triangleright Uniformly sample $n$ candidates
10:while $\{x^{\prime}\in Z:valid(x^{\prime})\land robust(x^{\prime})\}\neq\emptyset$ do
11:     $\eta\leftarrow 0.5\eta$
12:     $Z\sim_{n}Sphere(x^{cf},0,\eta)$
13:end while
14$\triangleright$ GrowingSpheres search: find closest, robust, and vaild CFEs
15:$a_{0}\leftarrow\eta$, $a_{1}\leftarrow 2\eta$
16:$Z\sim_{n}Sphere(x^{cf},a_{0},a_{1})$
17:while $\{x^{\prime}\in Z:valid(x^{\prime})\land robust(x^{\prime})\}==\emptyset$ do
18:     $a_{0}\leftarrow a_{1}$,
19:     $a_{1}\leftarrow a_{1}+\eta$
20:     $Z\sim_{n}Sphere(x^{cf},a_{0},a_{1})$
21:end while
22$\triangleright$ Return robust and valid CFE closest to the base CFE
23:return $\min_{x^{\prime}\in Z:valid(x^{\prime})\land robust(x^{\prime})}d(x^{\prime},x^{cf})$
@@
Proof sketch: In Theorem 2 $a+k$ represents the most optimistic situation, where all $k$ estimators agree that the counterfactual is robust, resulting in the most extreme right-skewed beta distribution and the highest attainable $\delta$ with $k$ estimators. A proof and detailed analysis of this formula can be found in App. C.